OpenStudy (anonymous):

Please help! A sculpture was purchased for $4,000. The value of the sculpture increases at a rate of 4% per year. After 10 years the sculpture will be worth $5,920.97. In how many years wil the sculpture be worth $10,000? Round to the nearest tenth of the year.

OpenStudy (amistre64):

power went out ....

OpenStudy (amistre64):

A = P (1+ r) ^t corresponds to: 10000 = 4000(1+.04)^t now we solve for t ln(A/P) ------ = t ln(1+r)

OpenStudy (anonymous):

? What does A, P r stand for again?

OpenStudy (amistre64):

i lined up the results so you could match them up :)

OpenStudy (amistre64):

They are just standard variables in the equations

OpenStudy (anonymous):

2.54 years?

OpenStudy (anonymous):

This one took me a while to remember how to solve. Use A = Pe^rt , P= 4000, A=10000, e will cancel with Ln : 10000 = 4000e^0.04(t) Divide both sides by 4000 to get rid of it... 2.5 = e^0.04(t) Ln(2.5) = Ln*e^0.04(t) t = Ln(2.5)/0.04 t = 22.9 years

OpenStudy (amistre64):

that was a good try :) but the Pert is for "continuous" compounding. Other than that, you did quite well ;)